OpenStudy (anonymous):

f(x)=(x-1)^3+2 describe the transformation

OpenStudy (nenadmatematika):

=x^3-3x^2+3x-1+2=x^3-3x^2+3x+1

OpenStudy (anonymous):

THANKS U :) hey i got a big problem if u dont mind helping its for a project and i am just stumped

OpenStudy (nenadmatematika):

well post it so I can see....:D

OpenStudy (anonymous):

A land developer owns a piece of land surrounding a lake. The developer needs to decide how to divide the land into lots for lakeside -homes and regular homes in order to make the maximum profit. Lake front home: The developerís cost is $150,000; the profit over cost is $50,000. Regular home: The developerís cost is $100,000; the profit over cost is $23,000. There is enough land for at most 20 homes. There must be a minimum of 10 regular homes. The developer can invest no more than $2,200,000. Use linear programming to find how many lake front and regular homes the developer should build. I need to know the constraints and the inequaltities and the objective function anddd if u could show me a graph lol :P if u do this UR MY HERO

OpenStudy (anonymous):

does it make any sense :(?

OpenStudy (nenadmatematika):

linear programming is not really my thing sorry :D I hope that someone will answer to you, it's more the field of programming and economics :D

OpenStudy (anonymous):

ok thank u buddy anyway for the other help :)